# Predictability of Performance in Communication Networks Under Markovian Dynamics

### v3

@@ III System Model and Problem Formulation
A third assumption for the remainder of this work is that the simplified posterior distribution mentioned above is stationary, meaning that
$\forall L,\Pr(Z_{n+L}\mid X_{n+L})=\Pr(Z_{n}\mid X_{n}).$ (4)
For ease of notation, we represent the probability mass function (pmf) of this distribution as $\Pr(Z_{n}=z\mid X_{n}=x)=r_{x}(z)$, where $z\in\mathcal{Z}$ and $x\in\mathcal{X}$.
As shown in Figure 2, this system forms an observable Markov model (OMM) where the sequence $\{X_{0},X_{1},...,X_{n}\}$ is a discrete-time, time-homogeneous Markov chain with finite state space $\mathcal{X}$. The transition probability from state $x$ to state $y$ is denoted by $P(x,y)=\Pr(X_{n+1}=y\mid X_{n}=x)$ which forms the transition probability matrix $P$. $L$ step state transition probability is dervied via $P^{L}(x,y)$. Moreover, we assume that $X$ is an irreducible and aperiodic Markov chain, therefore it has an stationary distribution $\pi$ that state probabilities converge to for large lead times. Then, for all $x,y$ we have $P^{L}(x,y)\rightarrow\pi(y)$ as $L\rightarrow\infty$. Also, we assume the chain is reversible, which is an assumption satisfied by all random walks on undirected graphs, as well as many other Markov chains such as queues:
@@
In this paper, we will examine more specialized Markov chains, particularly the Geo/Geo/1/K queue, which represents the discrete-time counterpart of the M/M/1/K queue. In this case, the packets arrive according to a Bernoulli process and the probability of arriving during a time slot is $\lambda=1-\bar{\lambda}$. Service times are geometrically distributed, and the probability of completion of a service during a slot is $\mu=1-\bar{\mu}$. The utilization factor is denoted by $\rho=\lambda/\mu$. The performance metric $Z_{n}$ will be the sojourn time of the queue, and the queue size will be the measurable state of the system $X_{n}$. As the number of states in our analysis should be finite, we consider a maximum length of $K$ for the queue.
To formulate predictability in general, we rely on the forecast and the marginal distributions. An event is considered unpredictable not because its distribution is entirely unknown, but because the forecast distribution doesn’t significantly differ from the marginal distribution. Based on that, we define predictability in the following definition.
@@
A system is unpredictable if the forecast and marginal distributions do not differ if
@@
According to this definition, the loss of predictability occurs when the future state $Z_{n+L}$ is statistically independent of the observations. Equivalently, we define the degree of predictability as total variation distance between the forecast and marginal distributions
@@
where the Total Variation distance (TV) is a statistical metric distance defined by
@@
for probability mass functions $p$ and $q$.
Total variation distance is a metric that ranges between zero and one and benefits from numerous properties for bounding and approximation. Intuitively, the system is unpredictable when the TV distance between the forecast and the marginal is very small, so the performance and observations are almost independent. On the contrary, the system is predictable when TV is close to 1 since the distributions are very different and that means the observations make a difference in our prediction of the performance. This framework implies that predictability is a combined property of the dynamical system and the observations. The observations might be irrelevant, leading to zero predictability, or the system conditions might change in a completely random manner, rendering the incorporation of the observations ineffective.
@@
This metric guarantees that if we have observations up to time $n$, then we will have informative forecasts up to time $n+L$, since the forecast distribution will always have minimum distance $\epsilon$ with the marginal distribution. Essentially, it reflects how quickly the observations lose their importance, as the system evolves quickly to its stationary state. Therefore, epsilon-predictable horizon can be essential in analyzing predictability of complex dynamic systems.
$X_{n-1}$ $X_{n}$ $X_{n+1}$ $Z_{n-1}$ $Z_{n}$ $Z_{n+1}$$\dots$
Figure 2: Dependencies in the performance model
@@ IV-B Markov Chain Spectral Properties Impact
This theorem suggests that the predictability upper bound decays geometrically, with the same rate that Markov chain converges to stationarity. Upper bounds in Equations 35 and 32 are different in only one term which is sensitive to the closeness of posterior distributions $r_{y}(z)$. The minimum value of $R$, i.e. $1$, occurs when all conditional distributions are similar and makes the predictability zero as was expected. On the other hand, we get the maximum, i.e. $K=|\mathcal{X}|$, when conditional distributions have minimal overlap.
@@ IV-C Predictability in Multi-Hop Systems
Obtaining predictability using the above method becomes less tractable as the dimensions of the system grow. In this context, identifying approximation or bounding techniques applicable to the problem is helpful. Below we propose a key lemma regarding the predictability in tandem systems where the end-to-end delay is subject to forecast.
@@
The predictability of a tandem multi-hop system, assuming independence of the hops, is upper-bounded via the sum of predictability of each hop as
@@ IV-D Geo/Geo/1/K Queue Predictability
In this subsection, we utilize the previously mentioned predictability concepts and methods to the Geo/Geo/1/K queue model. The target metric $Z_{n}$ for this forecast system is the sojourn time of the queue. The transition matrix in this queue is given by
$\Pr(X_{n}=j\mid X_{n-1}=i)=\\ \begin{cases}\lambda\bar{\mu}&\text{ if }j=i+1,0\leq i\leq K-1\\ \mu\bar{\lambda}&\text{ if }j=i-1,1\leq i\leq K\\ 1-\lambda\bar{\mu}&\text{ if }j=i=0\\ \lambda\mu+\bar{\lambda}\bar{\mu}&\text{ if }j=i,1\leq i\leq K-1\\ 1-\mu\bar{\lambda}&\text{ if }j=i=K\\ 0,&\text{ otherwise }\end{cases}$ (41)
@@
00 $1$dots $K$ $1-\lambda\bar{\mu}$$\lambda\bar{\mu}$$\lambda\mu+\bar{\lambda}\bar{\mu}$$\lambda\bar{\mu}$$\mu\bar{\lambda}$$\lambda\mu+\bar{\lambda}\bar{\mu}$$\lambda\bar{\mu}$$\mu\bar{\lambda}$$1-\mu\bar{\lambda}$$\mu\bar{\lambda}$
@@
$D_{n}(L)=\\ \frac{1}{2}\sum_{z}\biggl{|}\sum_{y=1}^{K}\frac{2}{K+1}\beta^{\frac{y-x}{2}}\sum_{k=1}^{K}\frac{\alpha_{k}^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \left[\sin\left(\frac{xk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(x+1)k\pi}{K+1}\right)\right]\\ \left[\sin\left(\frac{yk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(y+1)k\pi}{K+1}\right)\right]\text{NB}(z;y,\mu)\biggr{|},$ (42)
@@
$\alpha_{k}=\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\cos\left(\frac{k\pi}{K+1}\right)\text{ and }\beta=\frac{\lambda\bar{\mu}}{\mu\bar{\lambda}}.$
@@
$P^{L}(i,j)=\\ =\frac{1-\beta}{1-\beta^{K+1}}\beta^{j}+\frac{2}{K+1}\beta^{\frac{j-i}{2}}\sum_{k=1}^{K}\frac{\alpha_{k}^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \left[\sin\left(\frac{ik\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(i+1)k\pi}{K+1}\right)\right]\\ \left[\sin\left(\frac{jk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(j+1)k\pi}{K+1}\right)\right].$ (44)
When the queue is stable, i.e. $\mu>\lambda$, the stationary probabilities are known as
@@
Then, we derive predictability by replacing Equations 43, 44, and 45 in Theorem 1 to complete the proof.
@@
$\kappa=\frac{\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}{\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}.$ (47)
@@
In the following section we focus on numerical evaluations of the predictability analysis methods in different scenarios and configurations.
V Numerical Evaluations
@@
This section is dedicated to evaluating the proposed methods and examining the predictability in abstract systems in 3 subsections [^1] MATLAB code and reproducible results are available at https://github.com/samiemostafavi/predictability. First, we evaluate the spectral-based upperbound and approximations on predictability of the Geo/Geo/1/K queue. The second subsection is devoted to analyzing how dynamics of the queue affect its predictability and in the end multi-hop queues are analyzed when their observability is limited.
V-A Approximation and Upper Bound Validation

### v4

@@ III-C Markovian Model of System State
This system forms an observable Markov model (OMM) where the sequence $\{X_{0},X_{1},...,X_{n}\}$ is a discrete-time, time-homogeneous Markov chain with finite state space $\mathcal{X}$. The transition probability from state $x$ to state $y$ is denoted by $P(x,y)=\Pr(X_{n+1}=y\mid X_{n}=x)$ which forms the transition probability matrix $P$. $L$ step state transition probability is derived via $P^{L}(x,y)$. Moreover, we assume that $X$ is an irreducible and aperiodic Markov chain, therefore it has an stationary distribution $\pi$ that state probabilities converge to for large lead times. Then, for all $x,y$ we have $P^{L}(x,y)\rightarrow\pi(y)$ as $L\rightarrow\infty$. Also, we assume the chain is reversible, which is an assumption satisfied by all random walks on undirected graphs, as well as many other Markov chains such as queues:
@@
We will also examine more specialized Markov chains, particularly the Geo/Geo/1/K queue, which represents the discrete-time counterpart of the M/M/1/K queue. In this case, the packets arrive according to a Bernoulli process and the probability of arriving during a time slot is $\alpha=1-\bar{\alpha}$. Service times are geometrically distributed, and the probability of completion of a service during a slot is $\mu=1-\bar{\mu}$. The utilization factor is denoted by $\rho=\alpha/\mu$. The performance metric $Z_{n}$ will be the sojourn time of the queue, and the queue size will be the measurable state of the system $X_{n}$. As the number of states in our analysis should be finite, we consider a maximum length of $K$ for the queue which in actual implementations corresponds to the system’s buffer size. Accordingly, $K^{(m)}$ would denote the buffer size at hop $m$. Because the queue is of finite length, packets arriving when the buffer is full are dropped, making this a lossy system, underscoring the importance of considering the buffer size when analyzing packet delay or packet loss [47].
Furthermore, to explore the predictability of network performance in connected vehicles, we consider that these networks often exhibit variable channel quality (e.g., CQI) due to fluctuations in mobility, interference, and propagation conditions. While exact modeling of these variations can be highly complex, a random walk process provides a tractable approximation for the sequential dependency observed in the CQI. The CQI is a standard metric in cellular networks such as LTE and 5G, indicating link quality between a base station and the UE. CQI variations can be substantial or minimal, depending on factors such as mobility, and are important because the connection’s throughput depends on the CQI [48]. In this model, states are represented within a finite one-dimensional space of size $K$. At each step, the state either remains unchanged with a certain probability or transitions to a neighboring state with equal likelihood (except for at the reflecting boundary). The process maintains a uniform stationary distribution, meaning that for any state $y$, the probability $\pi(y)$ is consistently $1/K$. Each CQI level corresponds to a particular throughput distribution, and transition probabilities between CQI levels we estimate from real-world mobile vehicle connection quality measurements [48]. Random walk models offer a simplified yet effective way to represent the stochastic nature of vehicular movement. These models have had numerous applications in the study and analysis of vehicular mobility within connected vehicle environments [49, 50, 51].
III-D Predictability Model and Definitions
The primary objective of any performance predictor in this setting is to capture the relationship between the observations and future performance. The probability distribution of the performance metric, $L$ timeslots into the future, based on the observations collected up to the present moment, is the forecast distribution:
$\Pr(Z_{n+L}\mid O_{0:n}=o_{0:n}).$ (6)
The distribution of performance right after observations become available denoted by $\Pr(Z_{n}\mid O_{0:n})$ is the posterior distribution [52]. The marginal distribution of performance (or prior distribution) denoted by $\Pr(Z_{n})$ is related to the posterior distribution by the chain rule as it is obtained by integrating out all observations:
The marginal distribution constitutes a baseline for the information content about performance, disregarding all state observations. This distribution can be derived by randomly sampling performance values and remains stationary because: (i) the conditional distribution is assumed to be stationary (as stated in Equation 1), and (ii) the marginal system state distribution $\pi$ is also stationary. Consequently, the marginal performance distribution retains its stationarity. Stationarity of the marginal distribution implies that the overall long-term behavior of the system, in terms of performance, is consistent. While the system might experience temporary fluctuations due to different states and observations, the overall probability of experiencing various levels of performance remains the same over time. This is reasonable for systems that are designed to operate within certain performance bounds in the long run.
To formulate predictability, we rely on the forecast and the marginal distributions defined above. An event is considered unpredictable not because its distribution is entirely unknown, but because the forecast distribution does not significantly differ from the marginal distribution. Based on that, we define predictability in the following definition.
@@ III-D Predictability Model and Definitions
A system is unpredictable with lead time $L$ if the forecast and marginal distributions do not differ if
@@
According to this definition, the loss of predictability occurs when the future state $Z_{n+L}$ is statistically independent of the observations. Equivalently, we define the degree of predictability as the TV distance between the forecast and marginal distributions
@@
where the TV distance is a statistical metric distance defined by
@@
for probability mass functions $p$ and $q$. It is important to emphasize that this definition relies on the marginal distribution being stationary.
The TV distance is a metric that ranges between zero and one and benefits from numerous properties for bounding and approximation. Intuitively, the system is unpredictable when the TV distance between the forecast and the marginal is very small, so the performance and observations are almost independent. On the contrary, the system is predictable when TV is close to 1 since the distributions are very different and that means the observations make a difference in our prediction of the performance. This framework implies that predictability is a combined property of the dynamical system and the observations. The observations might be irrelevant, leading to zero predictability, or the system conditions might change in a completely random manner, rendering the incorporation of the observations ineffective.
@@
This metric guarantees that if we have observations up to time $n$, then we will have informative forecasts up to time $n+L$, since the forecast distribution will always have minimum distance $\epsilon$ with the marginal distribution. Essentially, it reflects how quickly the observations lose their importance, as the system evolves quickly to its stationary state. Therefore, the epsilon-predictable horizon is essential in analyzing the predictability of complex dynamic systems.
@@ IV-B Markov Chain Spectral Properties Impact
This theorem suggests that the predictability upper bound decays geometrically, with the same rate that Markov chain converges to stationarity. Upper bounds in Equations 36 and 33 are different in only one term which is sensitive to the closeness of posterior distributions $r_{y}(z)$. The minimum value of $R$, i.e. $1$, occurs when all conditional distributions are similar and makes the predictability zero as was expected. On the other hand, we get the maximum, i.e. $K=|\mathcal{X}|$, when conditional distributions have minimal overlap.
@@ IV-C Predictability in Multi-Hop Systems
Deriving predictability using the above method becomes less tractable as system complexity increases. In this context, identifying approximation or bounding techniques applicable to the problem is helpful. Below we propose a key lemma regarding the predictability in tandem systems where the end-to-end delay is subject to forecast.
@@
The predictability of a tandem multi-hop system, assuming independent hops, is upper-bounded by the sum of the predictability of each hop as follows:
@@ IV-D Geo/Geo/1/K Queue Predictability
In this subsection, we apply the previously described predictability concepts and methods to the Geo/Geo/1/K queue model, where $K$ denotes the maximum queue length or buffer capacity. We focus on the sojourn time (or packet delay) $Z_{n}$ as the primary performance metric for our forecasts. Finally, we offer an investigation on how predictability definitions extend to the case in which $Z_{n}$ represents the queue’s blocking to model packet loss ratio instead. The transition matrix in this queue is given by
$\Pr(X_{n}=j\mid X_{n-1}=i)=\\ \begin{cases}\alpha\bar{\mu},&\text{ if }j=i+1,0\leq i\leq K-1\\ \mu\bar{\alpha},&\text{ if }j=i-1,1\leq i\leq K\\ 1-\alpha\bar{\mu},&\text{ if }j=i=0\\ \alpha\mu+\bar{\alpha}\bar{\mu},&\text{ if }j=i,1\leq i\leq K-1\\ 1-\mu\bar{\alpha},&\text{ if }j=i=K\\ 0,&\text{ otherwise }\end{cases}$ (42)
@@
00 $1$dots $K$ $1-\alpha\bar{\mu}$$\alpha\bar{\mu}$$\alpha\mu+\bar{\alpha}\bar{\mu}$$\alpha\bar{\mu}$$\mu\bar{\alpha}$$\alpha\mu+\bar{\alpha}\bar{\mu}$$\alpha\bar{\mu}$$\mu\bar{\alpha}$$1-\mu\bar{\alpha}$$\mu\bar{\alpha}$
@@
$D_{n}(L)=\\ \frac{1}{2}\sum_{z}\biggl{|}\sum_{y=1}^{K}\frac{2}{K+1}\beta^{\frac{y-x}{2}}\sum_{k=1}^{K}\frac{\gamma_{k}^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \left[\sin\left(\frac{xk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(x+1)k\pi}{K+1}\right)\right]\\ \left[\sin\left(\frac{yk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(y+1)k\pi}{K+1}\right)\right]\text{NB}(z;y,\mu)\biggr{|},$ (43)
@@
$\gamma_{k}=\alpha\mu+\bar{\alpha}\bar{\mu}+2\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}\cos\left(\frac{k\pi}{K+1}\right)\text{ and }\beta=\frac{\alpha\bar{\mu}}{\mu\bar{\alpha}}.$
@@
$P^{L}(i,j)=\\ =\frac{1-\beta}{1-\beta^{K+1}}\beta^{j}+\frac{2}{K+1}\beta^{\frac{j-i}{2}}\sum_{k=1}^{K}\frac{\gamma_{k}^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \left[\sin\left(\frac{ik\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(i+1)k\pi}{K+1}\right)\right]\\ \left[\sin\left(\frac{jk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(j+1)k\pi}{K+1}\right)\right].$ (45)
When the queue is stable, i.e. $\mu>\alpha$, the stationary probabilities are known as
@@
Then, we derive predictability by replacing Equations 44, 45, and 46 in Theorem 1 to complete the proof.
@@
$\kappa=\frac{\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}}{\alpha\mu+\bar{\alpha}\bar{\mu}+2\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}}.$ (48)
@@
Finally, as noted previously, a Geo/Geo/1/K queue is a lossy system that may reject or drop arriving packets if the queue is full. We demonstrate that our proposed predictability framework can capture the packet loss ratio in this queueing scenario. Specifically, let $Z_{n}$ be an indicator variable describing whether an arriving packet at slot $n$ encounters a full queue, with $Z_{n}\in\{0,1\}$. Hence, $\Pr(Z_{n}=1)$ is the packet loss ratio at slot $n$, which we take as our performance metric in place of delay. In the following proposition, we provide an analytical derivation of the packet loss ratio’s predictability for a Geo/Geo/1/K queue.
Proposition 2
The predictability of blocking in a Geo/Geo/1/K queue for $L$ time slots in the future based on the observation of the queue size $x$ is given by
$D_{n}(L)=\biggr{|}\frac{2}{K+1}\beta^{\frac{K-x}{2}}\sum_{k=1}^{K}\frac{\gamma_{k}^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \left[\sin\left(\frac{xk\pi}{K+1}\right)-\sqrt{\beta}\sin\left(\frac{(x+1)k\pi}{K+1}\right)\right]\sin\left(\frac{Kk\pi}{K+1}\right)\biggr{|}.$ (49)
Proof:
Since $Z_{n}$ takes values in $\{0,1\}$, the definition of predictability reduces to
$D_{n}(L)=\bigl{|}\Pr(Z_{n}=1\mid X_{n}=x)-\Pr(Z_{n}=1)\bigr{|}.$ (50)
To determine the blocking probability, we apply the discrete‐time equivalent of PASTA for Bernoulli arrivals, which identifies the probability of blocking with the probability of the queue being in state $K$. Thus, when predicting $L$ steps ahead given $X_{n}=x$, the blocking probability is
$\Pr(Z_{n}=1\,\mid\,X_{n}=x)=P^{L}(x,K),$ (51)
and under steady‐state conditions, it is
$\Pr(Z_{n}=1)=\pi(K).$ (52)
Substituting these from Equations (45) and (46) into the predictability definition completes the proof. ∎
This proposition offers an analytic expression that quantifies how knowing the current queue size $x$ alters the predictability of near-future blocking in a Geo/Geo/1/K queue, compared to the steady-state behavior. The result underscores a non-trivial interplay among the buffer size $K$, the forecast horizon $L$, and the system state $x$. Even a relatively large $K$ can exhibit marked short-term predictability of blocking if the queue is already near capacity, whereas for a smaller $K$, small shifts in queue length can sharply change the likelihood of blocking. As $L$ increases, the system forgets its starting state due to Markov mixing, causing predictability to converge to zero. We will analyze this further in Section V.
V Numerical Results
@@
This section is dedicated to evaluating the proposed methods and examining the predictability in abstract systems in 3 subsections [^1] MATLAB code and results are accessible at https://github.com/samiemostafavi/predictability. First, we evaluate the connected vehicular scenario. Then, spectral-based upperbound and approximations on predictability of the Geo/Geo/1/K queue are analyzed. The last subsection is devoted to analyzing how dynamics of the queue affect its predictability and in the end multi-hop queues are analyzed when their observability is limited.
Throughout the section, our focus is on the relationship between predictability and lead time $L$. In our plots, lead time is consistently represented as the primary independent variable on the x-axis. A fundamental observation across all scenarios is that predictability decreases as the lead time $L$ increases. This reflects the inherent difficulty in predicting the future behavior of a system as we look further ahead. The rate of this decrease, which is often approximately exponential, is influenced by various system parameters, such as the maximum queue size $K$, the level of state aggregation, and the presence of bottlenecks in multi-hop systems. We will explore these influences in detail in the following subsections.
V-A Predictability Under the Random Walk Model
As a starting point, we employ a lazy random walk to capture the dynamics of vehicular connections with varying CQI in a cellular setting. In this scenario, the vehicle’s connection state—represented by $X_{n}$ for the CQI —follows a lazy random walk on a finite, one-dimensional state space of size $K$, with reflecting boundaries. Specifically, with probability $p$ the state remains unchanged, and with probability $\tfrac{1-p}{2}$ it transitions to each of the two adjacent states (reflecting at the boundaries). This random walk has a uniform stationary distribution, $\pi(y)=\tfrac{1}{K}$ for all $y$.
Each CQI state $X_{n}$ ($K=15$) modulates the network’s throughput, which we assume follows a Negative Binomial distribution $\mathrm{NB}(z;r,q)$. We set its parameters as
$r=0.105\cdot X_{n}+0.104,\quad q=-0.006\cdot X_{n}+0.135.$
These parameter values are derived from a real-world dataset in which LTE connection’s downlink throughput (in Mbps) and CQI were sampled at one-second intervals for both mobile (train, bus, car) and non-mobile scenarios in Cork city, Ireland [48]. See Appendix C for further details on how these parameters were derived.
From our empirical analysis, we observe that in vehicular connections the random walk remains in the same state with probability $p=0.6$, whereas in near-static scenarios $p=0.9$. This modeling approach thus links the random-walk-based CQI evolution directly to throughput distribution in mobile environments. Figure 4 illustrates the resulting posterior distributions given an observed connection CQI, along with the marginal distribution derived using Lemma 1, and a forecast distribution for 60 seconds after when CQI 15 was observed given the random walk model.
Figure 4: Posterior, marginal, and forecast distributions for a cellular connection’s downlink throughput, with CQI evolving under a random walk model with $p=0.6$ (mobile).
(a) Mobile vehicular connection with $p=0.6$
(b) Static connection with $p=0.9$
Figure 5: Predictability of a cellular connection’s downlink throughput, with CQI evolving under a random-walk model. The solid lines depict exact values at various lead times, while dotted and dashed lines represent two spectral upper bounds (1 and 2), respectively.
Figure 5 shows the exact predictability calculated using Theorem 1, along with two upper bounds described in Theorem 2: the full spectral-based upper bound in Equation 32 illustrated with dashed lines (UB1) and the spectral gap-based bound in Equation 33 (UB2) illustrated with dotted lines. Several key observations can be made from these results:
  1. 1.
    Decay patterns: All predictability curves show a decay as the lead time increases, reflecting the gradual loss of predictive power as we look further into the future. However, predictability remains higher for a longer duration in static connections (Figure 5(b)) because the measured CQI is more likely to stay constant, unlike the more variable behavior observed in mobile scenarios (Figure 5(a)).
  2. 2.
    State-dependent predictability: The first and last states ($x=1$ and $x=15$) offer higher predictability compared to middle states (e.g., $x=3$). This is because their corresponding conditional distributions differ more significantly from the marginal distribution.
  3. 3.
    Upper bound tightness: For smaller lead times $L$, the upper bounds show a greater distance from the true predictability values. In general, we observe that the upper bounds are tighter for predictability curves that decay more rapidly.
  4. 4.
    Spectral gap bound limitations: The spectral gap-based bounds do not differentiate between observed states $x$, as seen in Equation 36, and are typically more relaxed due to the additional simplifications were applied to them. This is due to the equiprobable stationary state probabilities of the Markov chain in this case.
These findings highlight the role of Markovian dynamics in shaping the predictability of network throughput, particularly in mobile environments where the state persistence probability influences forecasting accuracy. In the next subsection, we extend this analysis to queuing systems.
V-B Approximation and Upper Bound Validation